A building of 10m high.two objects are projected upwards the 10m high building then the other downwards,with the same unitial velocity of 20m/s at the same time.calculate the time they gona meet and the distances.
1
Expert's answer
2011-02-18T07:50:02-0500
Denote the distance from the ground where two objects should meet as X.
According to the equations of the motion for each object:
X = h - vt - gt2/2
X = 0 +vt - gt2/2

Thus 0 = h - 2vt, t = h/2v = 10/(2*20) = 0.25 [s]
X = 20*0.25 - 10*0.252/2 = 5 - 0.3125 = 4.6875 [m]